x+5/5=x-4/4
it is x plus 5 divided by 5 is equal to x minus 4 divided by 4.
I know the answer is 40, but i do not know how you work the problem to get that answer.
Thank you

Answer by mollukutti(30)

To make things look easier let us re-write the qs as
(x+5)/5 = (x-4)/4
or, 4(x+5)=5(x-4) By cross multiplication of the denominators
or, 4x + 20 = 5x -20
or, 4x - 5x = -20 -20
or, -x = - 40
or, x = 40 (elimination of negative signs)